OpenStudy (anonymous):

Solve each system of equations by graphing. y = 2x + 1 y = -x + 1

OpenStudy (anonymous):

y = 2x + 1 y = -x + 1 --------------------- 2x + 1 = -x + 1 solve for x

OpenStudy (anonymous):

I don't really know how to do that.

OpenStudy (anonymous):

Could you help me? I'm sorry :(

OpenStudy (anonymous):



OpenStudy (anonymous):

How did you do that all? I don't get it :(

OpenStudy (anonymous):

And how do I graph it?

OpenStudy (anonymous):

@some_someone

OpenStudy (anonymous):

Graph: y = 2x + 1 & Graph: y = -x + 1

OpenStudy (anonymous):

graph the two lines on a graph and then look for the point of intersection

OpenStudy (anonymous):

I don't get how you graph them?! :(

OpenStudy (anonymous):

You have to talk me through this, we learned it today and I was too scared to ask for help

OpenStudy (anonymous):

the line equation is in the form y=mx+b where m is the slope of the line, and b is the y-intercept

OpenStudy (anonymous):

Yes I get that part

OpenStudy (anonymous):

so for y = 2x + 1 identify the slope and the y-intercept

OpenStudy (anonymous):

remember: y = mx + b --where: m is the slope; & b is the y-intercept

OpenStudy (anonymous):

I did what the someone guy said and it was (0,1) but I think I need another one

OpenStudy (anonymous):

?

OpenStudy (anonymous):

graph y = 2x + 1

OpenStudy (anonymous):

How do you graph that?

OpenStudy (anonymous):

Could you do the graph for me as an example? I'm lost

OpenStudy (anonymous):

slope = rise / run = 2/1 b or y-intercept = 1

OpenStudy (anonymous):

What do I do with the (0,1) ?

OpenStudy (anonymous):

thats the answer but the question asking you get that by graphin

OpenStudy (anonymous):

Ohhhhh

OpenStudy (anonymous):

yeah (0,1) is the solution.

OpenStudy (anonymous):

Wait so I go up 2 and then to the right 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what about the other 1

OpenStudy (anonymous):

slope is -1 so you go down 1 and right 1

OpenStudy (anonymous):

I have to have 2 lines though

OpenStudy (anonymous):

everything clear now?

OpenStudy (anonymous):

I thoguht I was suppose to go up 2. On yours it says the point of intersection is 1

OpenStudy (anonymous):

ya thats where the two lines intersect

OpenStudy (anonymous):

could you use this and do it on here

OpenStudy (anonymous):

@dav21

OpenStudy (anonymous):

ok first im gonig to draw y = 2x + 1

OpenStudy (anonymous):

first plot the y-intercept which is at point (0,1) then go up 2 and right 1 cause thats the slope

OpenStudy (anonymous):

ok got that part

OpenStudy (anonymous):

ok connect all the dots and make a straight line

OpenStudy (anonymous):

ok got that part too. But I have to have another line and stuff. There has to be 2.

OpenStudy (anonymous):

ya now graph y = -x + 1

OpenStudy (anonymous):

I don't understand how you graph these unless you give me the points and tell me which way to go

OpenStudy (anonymous):

y = -x + 1, first plot the y-intercept

OpenStudy (anonymous):

the y-intercept is at 1

OpenStudy (anonymous):

same as the first line

OpenStudy (anonymous):

how many do I go to the right and up

OpenStudy (anonymous):

y = (-1)x + 1, because our slope is negative we go down 1 and right 1

OpenStudy (anonymous):

positive slope means go up and right, negative slope means go down and right

OpenStudy (anonymous):

Thanks! Can you help me with a few moer?

OpenStudy (anonymous):

@dav21

OpenStudy (anonymous):

x + y = 4 x + y = 2

OpenStudy (anonymous):

ok first lets get the two equations in y = mx + b form

OpenStudy (anonymous):

ok. I need help with that

OpenStudy (anonymous):

x + y = 4, bring the x on the right side

OpenStudy (anonymous):

ok now what

OpenStudy (anonymous):

do the same for the other line as well

OpenStudy (anonymous):

show me what you have so far

OpenStudy (anonymous):

I have x + y = 4

OpenStudy (anonymous):

x + y = 2

OpenStudy (anonymous):

x + y = 4, if we subtract x from both sides of the equation what do we have

OpenStudy (anonymous):

y = 4x?

OpenStudy (anonymous):